# UnrealZoo: Enriching Photo-realistic Virtual Worlds for Embodied AI

### v1

@@ D.1 Learning Curve
We provide the CQL loss curve under the 1 Env., 4 Envs. and 8 Envs. training setup. As shown in Figure 15, the offline model approaches convergence after two hours and we continued training for another three hours after nearing convergence, observing no significant further decrease in the loss. Note that the offline training was conducted on a Nvidia RTX 4090 GPU.
@@
Figure 16: The learning curves for RL-based navigation agent in two environments: Roof and Factory. We use A3C (Mnih et al., 2016) to learn the navigation policy via trial-and-error interactions. In the Factory (blue line plot), the number of asynchronous workers is set to 4, while in the Roof environment (orange line plot), the number of asynchronous workers is set to 6.
D.2 Evaluate Tracking Agents across 16 Unseen Environments
@@
D.3 Evaluate Tracking Agents across Unseen Social Environments
We select 4 environments from different categories as the testing environments, including StorageHouse, DesertRuins, TerrainDemo, and SurburNeighborhoodDay. We test the distraction robustness of the social tracking agents by adding different numbers of distractors (4, 8, 10) in the environment. The distractors randomly walk around the environment, which may produce various unexpected perturbations to the tracker, such as visual distractions, occlusion, or blocking the tracker’s path. As shown in Table 12, the tracking performance of the three agents steadily decays with the increasing number of distractors.

### v2

@@ D.1 Learning Curve
We provide the CQL loss curve under the 1 Env., 4 Envs. and 8 Envs. training setup. As shown in Figure 15, the offline model approaches convergence after two hours, and we continued training for another three hours after nearing convergence, observing no significant further decrease in the loss. Note that the offline training was conducted on an Nvidia RTX 4090 GPU.
@@
Figure 16: The learning curves for the RL-based navigation agent in two environments: Roof and Factory. We use A3C [27] to learn the navigation policy via trial-and-error interactions. In the Factory (blue line plot), the number of asynchronous workers is set to 4, while in the Roof environment (orange line plot), the number of asynchronous workers is set to 6.
D.2 Testing in 16 Unseen Environments
@@
D.3 Testing in Social Tracking Scenarios
We select 4 environments from different categories as the testing environments, including StorageHouse, DesertRuins, TerrainDemo, and SuburbNeighborhoodDay. We test the distraction robustness of the social tracking agents by adding different numbers of distractors (4, 8, 10) in the environment. The distractors randomly walk around the environment, which may produce various unexpected perturbations to the tracker, such as visual distractions, occlusion, or blocking the tracker’s path. As shown in Table 14, the tracking performance of the three agents steadily decays with the increasing number of distractors.
D.4 Cross-platform Evaluation
To evaluate the generalization ability of policies trained in UnrealZoo, we conducted cross-platform experiments in both simulation and real-world settings. During our investigation, we found that many existing simulation environments  [29, 52, 28] lack flexibility in task customization, making it difficult to add or modify task definitions. Among the available options, we selected ThreeDWorld [14] for its minimal migration overhead.
For evaluation, we used the publicly available ”Suburb Scene 2023” map and designated the robot Magnebot as the tracking target, as shown in Figure 17. We tested two policies trained under different environment settings: a single environment (1-Env) and eight diverse environments (8-Envs). Due to limitations in ThreeDWorld, specifically, the inability to retrieve the absolute position of Magnebot, we could not compute the Average Reward (AR) metric. Instead, we report the episode length (EL) and success rate (SR) as alternative indicators of tracking performance.
Beyond simulation, we further validated the trained policies in the real world, illustrated in Figure 17. In this scenario, the target person walks along an S-shaped path in an open area with various distractors. We conducted five trials for each policy. Since all trials were successful, the success rate alone could not reflect performance differences. Additionally, absolute position data was unavailable in the real-world setting, so we introduced two alternative evaluation metrics:
  • Average Deviation (AD): the average pixel offset between the target’s bounding box center and the image center.
  • IoU: the Intersection over Union between the current target bounding box and the initial bounding box.
These metrics assess the agent’s ability to maintain consistent and centered tracking under dynamic motion and environmental interference. The final results are listed in Table 15. Overall, the experimental findings highlight that training with diverse environments in UnrealZoo significantly improves the policy’s generalization ability and robustness.
Figure 17: From left to right: 1)We utilize ThreedWorld as the comparison simulator and use its built-in agent “Magnebot” as the target, evaluating the active tracking performance. 2) We deploy our trained policy in a wheel-based robot with distractors, evaluating the active tracking performance.
Table 15: Cross-platform evaluation on active visual tracking task
Platform ThreeDWorld Real World
EL/SR AD/IoU
1 Env. 377/0.65 764/328
8 Envs. 493/0.8 833/393
Appendix E Limitations and Discussions.
While our proposed environment provides diverse and complex scenarios for visual navigation, tracking, and other embodied vision tasks, it currently has some limitations that should be overcome in future work: 1) Limited Physical Simulation Fidelity. The current version focuses on visual rendering and interaction rather than the precision of physical interactions for control. We will consider enhancing physics engine integration for more accurate and realistic physical interactions. 2) Licensing Restrictions and Shareability. Due to marketplace content licensing restrictions, we are unable to open-source the complete environment source code for more in-depth customization. To address this, we will develop alternative scenes based on free assets or negotiate special licensing agreements with asset creators, and extend the UnrealCV command system to support more advanced customization in the released binary. 3) Limited Interactions. Interaction capabilities are primarily limited to pre-defined objects, and multi-agent interactions remain relatively basic, lacking complex social dynamics. We will extend to more complex interaction tasks such as object manipulation, tool use, and environmental modification, and implement more sophisticated social agent behavior models, including dialogue and emotional expression.